# Transport properties through alternating borophene and graphene superlattices

### v1

@@ IV.2 Multiple barrier ($n$ cells)
To show how the number of cells $n$ affects the transport properties, we plot transmission as a function of energy for different values of $n$ in Fig. 7. For $n=1$, we observe that the transmission is almost perfect, with a slight difference between the two junctions. The transmission shows a consistent pattern as $n$ increases, suggesting resilience to fluctuations in $n$.
We see that, despite increasing $n=20,30$, the positions and numbers of resonant peaks indicate that the BGB junction retains complete transparency even at an incidence angle of $\theta=10^{\circ}$. Crucially, this transparency remains unaffected by variations in well widths. Resonances of the same kind are found for both junctions, with resonances in BGB being more significant than in GBG. To compare with our results, it is worth noting that at normal incidence, a structure resembling monolayer graphene consistently remains perfectly transparent [46]. By further increasing $n$, we observe that at $n=30$, band gaps begin to emerge. Indeed, the BGB junction exhibits three gaps, whereas the GBG junction has two. All these gaps fall within the same energy range, as depicted in the band structure shown in Fig. 2.
In Fig. 8, we plot transmission as a function of the energy and incident angle in BGB (a) and GBG (b) junctions through $n=20$ barriers, with $V_{0}=50$ meV and $d=w=20$ nm. We find that the properties of graphene, which are widely known, differ significantly from those of borophene-graphene superlattices. Unlike graphene, where the directions of the wave vector and the group velocity are collinear, our results in BGB and GBG show unique patterns in how electrons cross barriers, depending on both energy $E$ and angle $\theta$. Anisotropic dispersion in borophene superlattices introduces noncollinear directions and gives access to valley birefringence [36], whereas Klein tunneling in graphene is valley-degenerate and occurs mainly at normal incidence [42]. Our results indicate that, in contrast to graphene, where certain energies lead to a complete blocking of electron motion, in the borophene arrangement, certain energies allow electrons to pass through barriers more freely. This effect is particularly pronounced in the GBG junction for energies above $80$ meV, while the behavior of the BGB junction is different. Note that for the BGB junction, the appearance of resonances in the transmission at $E=80$ meV and $E=136$ meV occurs only when the incident angle is negative. This is in contrast to the results obtained in [36, 42], where the resonances appear only for $E<20$ meV. Under the same condition and at $E=50$ meV, the transmission differs from zero for the BGB junction.
@@
Fig. 10 illustrates how the transmission probability varies with respect to the barrier width $d$ for BGB (a) and GBG (b) junctions across multiple barriers. As $d$ increases, we observe two distinct transmission gaps whose widths progressively increase with increasing the barrier height $V_{0}$. Additionally, Fig. 10 demonstrates the appearance and disappearance of transmission gaps across different energy regions as $d$ increases.
Fig. 11 shows the transmission probability as a function of well width $w$. At $E=50$ meV, for $w=0$, we find the emergence of a transmission gap corresponding to a single barrier. Remarkably, as $d$ increases, this gap gets smaller while other transmission gaps arise. As $w$ rises further, these gaps become transmission gaps, first growing and then shrinking again. This phenomenon raises the prospect of using well width increases to create electron wave channels that are thinner [42]. Moreover, Fig. 11 illustrates how the center of each transmission gap gradually shifts to the left as $d$ rises, moving neighboring gaps closer together.
@@
Figure 14: (Color online) Ballistic conductivity and Fano factor at the Dirac point as a function of the barrier height $V_{0}$ for BGB (a) and GBG (b), with $E=0$, $V_{\infty}=-\infty$ and $D/d=5$.
@@
At this level, we discuss the effect of tilted energy on conductivity $\sigma=Gd/D$ and Fano factor $F$ in the ballistic regime. Such ballistic transport requires heavy doping of regions 1 and 3 with the electrostatic potential $V(x)=V_{\infty}$ for $x<0$ and $x>d$ and $V(x)=V_{0}$ for $0<x<d$. In Fig. 14, we present the ballistic conductivity and and Fano factor as functions of the barrier height $V_{0}$. At the Dirac point ($V_{0}=0$) and for $V_{\infty}=-\infty$, for the BGB junction in Fig. 14 (a), we observe the universal minimum ballistic conductivity $G_{0}/\pi$ and the Fano factor $F=0.333$, similarly to SLG [49]. Conversely, for the GBG junction in Fig. 14 (b), we have the results $1.295G_{0}/\pi$ and $F=0.343$. Notably, in GBG junctions, the conductivity is larger and the Fano factor smaller for $V_{0}<0$ compared to $V_{0}>0$. Additionally, it is noteworthy that BGB junction exhibits a symmetry at the Dirac point, in contrast to GBG junction.
Fig. 15 shows the ballistic conductivity and Fano factor  as a function of the width over length ratio $D/d$. For the BGB junction in Fig. 15 (a), at the Dirac point, an interesting behavior is noticed as the limiting case $D/d\rightarrow\infty$. Specifically, we note that the ballistic conductivity $\sigma$ tends towards $G_{0}/\pi$ and the Fano factor $F$ converges to $0.333$, as reported in [49]. Conversely, for the GBG junction in Fig. 15 (b), we see that $\sigma$ approaches $1.295G_{0}/\pi$, and the $F$ converges to $0.343$ when $D/d\rightarrow\infty$.  These results shed light on the relation between the length ratio of the junction and its ballistic transport, offering insights into the underlying electronic properties of the BGB and GBG junctions.
@@
We have studied the transport properties of BGB and GBG junctions by separately treating the cases of a single barrier and multiple barriers in a superlattice configuration. In the framework of the first case, we have demonstrated that the two junctions have exceptional transport properties, particularly showing perfect transmission at normal incidence. We have seen that a decrease in the barrier height leads to a loss of transmission. By varying the energy and transverse wave vector, we observed pronounced scattered transmission at normal incidence. In the case of multiple barriers, we found a strong superlattice transmission, with the BGB junction retaining perfect transparency even at non-zero incident angles. It turns out that by varying the incident energy, many gaps appear in the transmission probability, and by adjusting the number of cells, incident angle, and barrier parameters, one can modify the number, width, and position of these transmission gaps.
Under diffuse transport conditions, significant fluctuations in conductance and Fano factor were observed, highlighting the high sensitivity of BGB and GBG junctions to changes in physical parameters. Notably, distinct patterns emerged when comparing the behavior of BGB and GBG junctions, particularly with respect to their responses to changes in barrier height. In ballistic transport, our findings suggest that the minimum conductivity coincides with a maximum Fano factor. Moreover, we have noticed significant differences between BGB and GBG junctions, especially for conductivity and the Fano factor, which are influenced by the variation of the barrier height. Furthermore, our study of the length ratio (geometrical factor) has revealed intriguing patterns, as conductivity and the Fano factor approach specific values as the ratio approaches infinity. These results could offer valuable insights for developing electronic junctions utilizing graphene and borophene materials.

### v2

@@ IV.2 Multiple barrier ($n$ cells)
To show how the number of cells $n$ affects the transport properties, we plot transmission as a function of energy for different values of $n$ in Fig. 7. For $n=1$, we observe that the transmission is almost perfect, with a slight difference between the two junctions. The transmission shows a consistent pattern as $n$ increases, suggesting resilience to fluctuations in $n$. Our results are in good agreement with those developed in other studies on graphene subjected to multiple electrostatic barriers [53, 54, 55]. We see that, despite increasing $n=5,20,30$, the positions and numbers of resonant peaks indicate that the BGB junction retains complete transparency even at an incidence angle of $\theta=10^{\circ}$. Crucially, this transparency remains unaffected by variations in well widths. Resonances of the same kind are found for both junctions, with resonances in BGB being more prominent than in GBG. To compare with our results, it is worth noting that at normal incidence, a structure resembling monolayer graphene consistently remains perfectly transparent [56]. At $n=30$, we observe that transmission gaps begin to emerge. Indeed, the BGB junction exhibits three gaps, whereas the GBG junction has two. All these gaps fall within the same energy range, as depicted in the band structure shown in Fig. 2.
In Fig. 8, we plot transmission as a function of the energy and incident angle in BGB (a) and GBG (b) junctions through $n=20$ barriers, with $V_{0}=50$ meV and $d=w=20$ nm. We find that the properties of graphene, which are widely known, differ drastically from those of borophene-graphene superlattices. Unlike graphene, where the directions of the wave vector and the group velocity are collinear, our results in BGB and GBG show unique patterns in how electrons cross barriers, depending on both energy $E$ and angle $\theta$. Anisotropic dispersion in borophene superlattices introduces noncollinear directions and gives access to valley birefringence [36], whereas Klein tunneling in graphene is valley-degenerate and occurs mainly at normal incidence [46]. Our results indicate that, in contrast to graphene, where certain energies lead to a complete blocking of electron motion, in the borophene arrangement, certain energies allow electrons to pass through barriers more freely. This effect is particularly pronounced in the GBG junction for energies above $80$ meV, while the behavior of the BGB junction is different. Note that for the BGB junction, the appearance of resonances in the transmission at $E=80$ meV and $E=136$ meV occurs only when the incident angle is negative. This is in contrast to the results obtained in [36, 46], where the resonances appear only for $E<20$ meV. Under the same condition and at $E=50$ meV, the transmission differs from zero for the BGB junction.
@@
Fig. 10 illustrates how the transmission probability varies with respect to the barrier width $d$ for BGB (a) and GBG (b) junctions across multiple barriers. As $d$ increases, we observe two distinct transmission gaps whose widths progressively increase with increasing the barrier width $d$. Additionally, Fig. 10 demonstrates the appearance and disappearance of transmission gaps across different energy regions as $d$ increases.
Fig. 11 shows the transmission probability as a function of well width $w$. At $E=50$ meV, for $w=0$, we find the emergence of a transmission gap corresponding to a single barrier. Remarkably, as $w$ increases, this gap gets smaller while other transmission gaps arise. As $w$ rises further, these gaps become transmission gaps, first growing and then shrinking again. This phenomenon raises the prospect of using well width increases to create electron wave channels that are thinner [46]. Moreover, Fig. 11 illustrates how the center of each transmission gap gradually shifts to the left as $d$ rises, moving neighboring gaps closer together.
@@
Figure 14: (Color online) Ballistic scaled conductance and Fano factor at the Dirac point as a function of the barrier height $V_{0}$ for BGB (a) and GBG (b), with $E=0$, $V_{\infty}=-\infty$ and $D/d=5$.
@@
At this level, we discuss the effect of tilted energy on scaled conductance $\sigma=Gd/D$ and Fano factor $F$ in the ballistic regime. Such ballistic transport requires heavy doping of regions 1 and 3 with the electrostatic potential $V(x)=V_{\infty}$ for $x<0$ and $x>d$ and $V(x)=V_{0}$ for $0<x<d$. In Fig. 14, we present the ballistic scaled conductance and and Fano factor as functions of the barrier height $V_{0}$. At the Dirac point ($E=V_{0}=0$) and for $V_{\infty}=-\infty$, for the BGB junction in Fig. 14 (a), we observe the universal minimum ballistic scaled conductance $G_{0}/\pi$ and the Fano factor $F=0.333$, similarly to SLG [50]. Conversely, for the GBG junction in Fig. 14 (b), we have the results $1.295G_{0}/\pi$ and $F=0.343$. Notably, in GBG junctions, the scaled conductance is larger and the Fano factor smaller for $V_{0}<0$ compared to $V_{0}>0$. Additionally, it is noteworthy that BGB junction exhibits a symmetry at the Dirac point, in contrast to GBG junction.
Fig. 15 shows the ballistic scaled conductance and Fano factor as a function of the width over length ratio $D/d$. For the BGB junction in Fig. 15 (a), at the Dirac point ($E=V_{0}=0$), an interesting behavior is noticed as the limiting case $D/d\rightarrow\infty$. Specifically, we note that the ballistic scaled conductance $\sigma$ tends towards $G_{0}/\pi$ and the Fano factor $F$ converges to $0.333$, as reported in [50]. Conversely, for the GBG junction in Fig. 15 (b), we see that $\sigma$ approaches $1.295G_{0}/\pi$, and the $F$ converges to $0.343$ when $D/d\rightarrow\infty$. These results shed light on the relation between the length ratio of the junction and its ballistic transport, offering insights into the underlying electronic properties of the BGB and GBG junctions.
@@
We have studied the transport properties of BGB and GBG junctions by separately treating the cases of a single barrier and multiple barriers in a superlattice configuration. In the framework of the first case, we have demonstrated that the two junctions have exceptional transport properties, particularly showing perfect transmission at normal incidence. We have seen that a decrease in the barrier height leads to a loss of transmission. By varying the energy and transverse wave vector, we observed pronounced scattered transmission at normal incidence. In the case of multiple barriers, we found a strong superlattice transmission, with the BGB junction retaining perfect transparency even at non-zero incident angles. It turns out that by varying the incident energy, many gaps appear in the transmission probability, and by adjusting the number of cells, incident angle, and barrier parameters, one can modify the number, width, and position of these transmission gaps. Strong oscillations in conductance and Fano factor were observed under diffuse transport conditions, demonstrating the extreme sensitivity of BGB and GBG junctions to variations in physical parameters. Notably, distinct patterns emerged when comparing the behavior of BGB and GBG junctions, particularly with respect to their responses to changes in barrier height. In ballistic transport, our findings suggest that the minimum scaled conductance coincides with a maximum Fano factor. We also observed considerable differences between BGB and GBG junctions, especially for scaled conductance and the Fano factor, which are influenced by the variation of the barrier height. Furthermore, our study of the length ratio (geometrical factor) has revealed intriguing patterns, as scaled conductance and the Fano factor approach specific values as the ratio approaches infinity. These results could offer valuable insights for developing electronic junctions utilizing graphene and borophene materials.